### v1

# LayeredMAPF: a decomposition of MAPF instance without compromising solvability

Abstract
Generally, the calculation and memory space required for multi-agent path finding (MAPF) grows exponentially as the number of agents increases. This often results in some MAPF instances being unsolvable under limited computational resources and memory space, thereby limiting the application of MAPF in complex scenarios. Hence, we propose a decomposition approach for MAPF instances, which breaks down instances involving a large number of agents into multiple isolated subproblems involving fewer agents. Moreover, we present a framework to enable general MAPF algorithms to solve each subproblem independently and merge their solutions into one conflict-free final solution, without compromising on solvability. Unlike existing works that propose isolated methods aimed at reducing the time cost of MAPF, our method is applicable to all MAPF methods. In our results, we apply decomposition to multiple state-of-the-art MAPF methods using a classic MAPF benchmark [^1] https://movingai.com/benchmarks/mapf.html. The decomposition of MAPF instances is completed on average within 1s, and its application to seven MAPF methods reduces the memory usage and time cost significantly, particularly for serial methods. To facilitate further research within the community, we have made the source code of the proposed algorithm publicly available [^2] https://github.com/JoeYao-bit/LayeredMAPF.
Index Terms:
Multi-Agent Path Finding, graph theory.
I Introduction
Multi-agent pathfinding (MAPF), as its name suggests, computes a set of collision-free paths for multiple agents from their respective starting locations to target locations.
Existing methods for MAPF are capable of determining optimal or bounded suboptimal solutions, but efficiency remains a key factor limiting its application. Researchers have proposed novel methods to address this issue, such as trading off solution quality to reduce runtime, reducing search branch factors, and solving agents independently based on their priorities. However, these techniques are often not applicable to all MAPF methods, as they are limited to specific types of MAPF problems.
Motivated by the phenomenon that the cost of solving MAPF instances grows exponentially as the number of agents increases [li2022mapf], we propose a novel approach to reduce the cost of MAPF methods by decomposing a MAPF instance into multiple smaller subproblems. These subproblems are solved independently, while considering the solutions of other subproblems as dynamic obstacles.
This idea bears resemblance to Priority-Based Search (PBS) [ma2019searching], which assigns a unique priority to each agent and solves agents separately in decreasing priority order. PBS can be viewed as decomposing a MAPF instance into subproblems, each involving only one agent. While PBS is efficient, it lacks a guarantee of completeness. In contrast, our approach aims to decompose MAPF instances without compromising their solvability.
We formulate the decomposition of MAPF instances as a progressive optimization problem. Initially, we decompose the MAPF instance into multiple subproblems without restricting the order of solving, and then further split them into smaller subproblems with a limited solving order. To ensure solvability and minimize the size of each decomposition, we evaluate the solvability of each step of decomposition, allowing only those decompositions that do not compromise solvability. As a result, we demonstrate the performance of our method across various maps and illustrate its improvement on multiple cutting-edge MAPF methods.
Figure 1: These figures illustrate the complete process of decomposing a MAPF instance, which includes decomposing agents into initial clusters (a type of subproblem), further decomposing initial clusters into smaller clusters, and decomposing clusters into levels (another type of subproblem). For more details, please refer to Section IV.
The contributions of this manuscript are listed as follows:
1. We introduce a novel problem of decomposing a MAPF instance into solvable subproblems to reduce the solving cost. Additionally, we propose a novel method that decomposes a MAPF instance into multiple subproblems without compromising solvability. An overview of our method is depicted in Fig. 1.
2. We present a framework that enables general MAPF methods to solve subproblems independently and merge their solutions to obtain the solution of the original problem.
3. We conduct extensive testing to evaluate how the decomposition of MAPF instances influences various MAPF methods using classic MAPF datasets. We evaluate the impact in terms of time cost, memory usage, success rate, and solution quality.
The remainder of this article is organized as follows. In Section II, we review related problems and methods aimed at reducing the solving cost of MAPF instances. We then define basic terminology in Section III and introduce our method in Section IV. Our test results regarding the performance of decomposition under various maps are presented in Section V, followed by an examination of how decomposition benefits multiple cutting-edge MAPF methods in Section VI. Finally, Section VII concludes this article.
II Related works
In this section, we provide an overview of existing works aimed at reducing the time cost of MAPF methods. MAPF methods typically involve searching for paths connecting the starting and target locations for multiple agents. Broadly, there are two approaches to determining the paths of agents:
1, move one agent from its start to target while keeping other agents static and attempting to avoid conflicts with the paths of other agents. This approach is adopted by CBS-based methods, HCA*, LNS, and PBS.
2, move all agents simultaneously by avoiding conflicts with each other at each time step. This approach is employed by LaCAM, PIBT, and Push and Swap.
We refer to the first approach as serial MAPF methods and the second approach as parallel MAPF methods. This distinction is crucial in understanding how the decomposition of a MAPF instance applies to them. Further details about this difference can be found in Section IV-D.
II-A Conflict based search
Conflict-Based Search (CBS) [sharon2015conflict] is a two-level complete and optimal multi-agent pathfinding algorithm. At the high level, CBS employs a conflict tree in which nodes represent conflicts between paths of two agents. By splitting a conflict, CBS obtains constraints for the related two agents, respectively. The low level utilizes heuristic single-agent searches to find new paths that satisfy all constraints imposed by the high-level conflict tree node. Therefore, CBS operates as a serial MAPF method. If CBS finds a new path, it checks for conflicts between the new path and previous paths. If no new conflicts are found, CBS exits with a conflict-free solution; otherwise, it inserts the new conflict as nodes into the conflict tree and repeats the process of splitting conflicts and searching for new paths.
While CBS can find optimal solutions, it can be time-consuming in certain cases. To accelerate CBS, various improvements have been proposed.
II-A1 Trade off solution quality for efficency
Bounded CBS (BCBS) [hatem2013bounded] incorporates focal search into the low-level search of CBS, enabling the low-level search to consider avoiding conflicts with other agents’ paths and generate bounded suboptimal solutions. Enhanced CBS (ECBS) [hatem2013bounded] utilizes the same low-level search as BCBS and applies focal search to the high-level search of CBS, aiming to minimize the number of nodes from the current conflict tree (CT) node to the CT node representing the conflict-free solution. Similar to BCBS, ECBS also produces bounded suboptimal solutions.
Explicit Estimation CBS (EECBS) [li2021eecbs] employs online learning to estimate the cost of the solution beneath each high-level node. It utilizes Explicit Estimation Search (EES) [thayer2011bounded] to select high-level nodes for expansion, mitigating two drawbacks of ECBS: (1) the cost of the solution beneath a CT node $N$ might exceed the cost of $N$ and thus could surpass the suboptimality bound; (2) the lower bound of ECBS seldom increases, leading to difficulties in finding a solution within a reasonable time if the optimal sum of costs is not within the initial suboptimality bound.
II-A2 Bypassing conflicts
Bypassing Conflicts [boyarski2015don] is a conflict resolution technique that alters the paths of agents involved in a chosen conflict instead of splitting a conflict tree (CT) node and searching for paths to avoid related constraints. When expanding a CT node $N$ and generating child CT nodes, if the cost of a child CT node equals the cost of $N$ and the number of conflicts of the paths in the child node is smaller than the number of conflicts of $N$, then the child node is used to replace $N$, and all generated child CT nodes are discarded. Otherwise, $N$ is split as before. It has been demonstrated that bypassing conflicts often results in smaller CTs and reduces the runtime of CBS.
II-A3 Prioritizing conflicts
Prioritizing Conflicts [boyarski2015don] is a conflict-selection technique. A conflict is considered cardinal if, when CBS uses the conflict to split a conflict tree (CT) node $N$, the costs of both child CT nodes are larger than the cost of $N$. It is semi-cardinal if the cost of one child CT node is larger than the cost of $N$, and the cost of the other node is not. It is non-cardinal if the costs of both child CT nodes are equal to the cost of $N$. CBS can significantly enhance its efficiency by resolving cardinal conflicts first, then semi-cardinal conflicts, and finally non-cardinal conflicts. This prioritization is effective because generating CT nodes with larger costs first typically improves the lower bound of the CT (i.e., the minimum cost of the CT nodes in the open list) faster, thereby producing a smaller conflict tree and accelerating CBS.
II-A4 Symmetry reasoning
S
Symmetry Reasoning [li2019symmetry, li2020new] is a technique aimed at avoiding the repeated resolution of conflicts between the same pair of agents due to symmetric paths and conflicts. It efficiently identifies each symmetry and resolves it through a single splitting action with specialized constraints. This process results in a smaller conflict tree and reduces the time cost of CBS.
II-A5 Weighted Dependency Graph (WDG)
The Weighted Dependency Graph (WDG) heuristic [li2019improved] is an admissible heuristic employed in the high-level search of CBS. It operates by constructing a weighted dependency graph for each conflict tree (CT) node $N$. In these graphs, vertices represent agents, and edges denote dependency between agents. This dependency is defined as follows: the minimum sum of costs of their conflict-free paths satisfying $N$’s constraints (computed via solving a 2-agent MAPF instance using CBS) is greater than the sum of costs of their paths in $N$’s paths (the shortest paths satisfying $N$’s constraints but not necessarily conflict-free). The edge weights signify the difference between the minimum sum of costs of conflict-free paths satisfying $N$’s constraints and the sum of costs of their paths in $N$’s paths.
The value of the edge-weighted minimum vertex cover of the graph serves as an admissible heuristic for the high-level search of CBS. Despite the runtime overhead associated with building the weighted dependency graphs and determining their edge-weighted minimum vertex cover, incorporating the WDG heuristic often leads to a reduction in conflict tree size and a decrease in the runtime of CBS.
II-B Large neighborhood search
At one end of the spectrum are (bounded-sub) optimal algorithms capable of finding high-quality solutions for small problems. At the opposite end are unbounded-suboptimal algorithms, which can handle large problems but typically produce low-quality solutions. Li et al. proposed MAPF-LNS [li2021anytime], presenting a third approach that combines the strengths of both.
MAPF-LNS starts with a given solution and deletes a portion of it, referred to as a neighborhood, while treating the remaining part as fixed. This simplifies the original problem, making it easier to solve. If the newly found solution is superior to the current one, it replaces the existing solution. Consequently, MAPF-LNS is both complete and capable of iteratively enhancing solution quality towards near-optimality. During each iteration, LNS selects a subset of agents and updates their paths without modifying those of other agents. Thus, MAPF-LNS operates as a serial MAPF method. MAPF-LNS can employ any desired approach to solve the simplified problem, provided it can account for the fixed information.
Building on this work, Li et al. introduced MAPF-LNS2 [li2022mapf], which efficiently finds a solution (rather than improving a given solution) for a MAPF instance. Initially, MAPF-LNS2 invokes a MAPF algorithm to solve the instance and acquire a (partial or complete) plan. For agents lacking a path, MAPF-LNS2 devises a path that minimizes collisions with existing paths.
II-C Priority based search
Priority-Based Search (PBS) [ma2019searching] is an incomplete, suboptimal algorithm designed for prioritized planning. It employs a depth-first search at the high level to dynamically construct a priority ordering, thereby forming a priority tree (PT). When confronted with a collision, PBS greedily determines which agent should be assigned a higher priority. It efficiently backtracks and explores alternative branches only if no solution is found in the current branch. Consequently, PBS incrementally constructs a single partial priority ordering until all collisions are resolved.
Once each agent is assigned a unique priority, PBS computes a minimum-cost path (in priority order) from its starting vertex to its target vertex without colliding with the paths of agents with higher priorities that have already been planned. Therefore, PBS operates as a serial MAPF method.
PBS ranks among the most efficient methods for solving MAPF. However, prioritized planning with an arbitrary priority ordering does not guarantee completeness or optimality in general [ma2019searching].
II-C1 Greedy PBS
Given that PBS becomes less effective for MAPF instances with high densities of agents and obstacles, [chan2023greedy] introduced Greedy PBS (GPBS). GPBS utilizes greedy strategies to enhance PBS by minimizing the number of collisions between agents. In essence, GPBS employs the numbers of conflicts and conflicting pairs as heuristics to guide the search on its low and high levels, respectively.
PBS can be regarded as a special decomposition of MAPF instances, where each subproblem involves only one agent. Unlike PBS, our decomposition does not necessitate that every subproblem contain only one agent, while still preserving completeness. Additionally, our decomposition serves as an auxiliary component rather than an isolated method, making it applicable to a wide range of MAPF algorithms.
II-D PIBT and LaCAM
II-D1 Priority Inheritance with Backtracking (PIBT)
PIBT [okumura2019priority] is an incomplete and suboptimal MAPF method that assigns a unique priority to each agent at every timestep. This prioritization ensures that all movements are ordered, and all agents act within a single timestep according to their priority, making it a parallel MAPF method. Priority inheritance is utilized to effectively handle priority inversion in path adjustment within a small time window, and a backtracking protocol prevents agents from becoming stuck.
Based on priorities and assuming only local interactions, PIBT is easily adaptable to decentralized contexts. Intuitively, PIBT is well-suited for large environments where intensive communication is prevalent, and path efficiency improves with a lower density of agents.
In PIBT, for each timestep, each agent must evaluate distances from surrounding nodes to its goal. While this operation could be implemented by calling A* on demand, it may also become a bottleneck. To address this issue, the authors of PIBT proposed PIBT+ [okumura2022priority], which saves computation time by preparing distance tables from each agent’s goal.
II-D2 Lazy constraints addition for MAPF (LaCAM)
LaCAM [okumura2023lacam] is a complete and suboptimal two-level MAPF method. At the high level, it explores a sequence of configurations, where each search node corresponds to one configuration. At the low level, it searches for constraints specifying which agents go where in the next configuration, making it a parallel MAPF method.
For each high-level node, LaCAM performs a low-level search that generates constraints determining the agents’ positions in the next configuration. Successors at the high level (i.e., configurations) are lazily generated while adhering to constraints from the low level, resulting in a significant reduction in search effort. The successor generator of LaCAM is implemented using PIBT.
It is noteworthy that when LaCAM encounters a known configuration, it reinserts the corresponding high-level node into the open set (which stores nodes for high-level search). This action prevents repeated appearance of configurations, thereby preventing LaCAM from accepting external paths as dynamic obstacles to avoid, as all agents may need to wait at times to avoid conflicts with external paths.
Okumura et al. proposed LaCAM2 [okumura2023improving], which introduces two enhancements to LaCAM. The first enhancement is its anytime version, called LaCAM*, which converges to optimality eventually, provided that solution costs accumulate transition costs. The second enhancement adds a swap action to the successor generator of LaCAM, enabling the quick generation of initial solutions.
III Preliminaries
III-A Basic definitions
In this section, we present fundamental definitions of multi-agent pathfinding. Since our method is dimension-independent, the provided definitions apply to any-dimensional grid space $\mathcal{C}_{\mathcal{N}}$.
III-A1 Grid space
Let $\mathcal{C}_{\mathcal{N}}$ ($\mathcal{N}\geq 2$) denote a finite $\mathcal{N}$-dimensional integer Euclidean space, where the size of the space is defined as $\mathcal{D}\rightarrow\mathcal{C}_{\mathcal{N}}$, with $\mathcal{D}=\{d_{1},d_{2},\ldots,d_{i},\ldots,d_{\mathcal{N}}\}$ and $d_{i}\in\mathbb{N}$. The coordinates of an element g in this space are defined as a vector $(x_{1},x_{2},\ldots,x_{i},\ldots,x_{\mathcal{N}})$, where $x_{i}\in([0,d_{i})\cap\mathbb{N})$.
III-A2 Grid states
There are only two possible states for a grid/element in $\mathcal{C}_{\mathcal{N}}$: passable or unpassable. The set of all passable grids in $\mathcal{C}_{\mathcal{N}}$ is denoted as $\mathcal{F}\rightarrow\mathcal{C}_{\mathcal{N}}$, while the set of all unpassable grids is denoted as $\mathcal{O}\rightarrow\mathcal{C}_{\mathcal{N}}$. Therefore, $(\mathcal{F}\rightarrow\mathcal{C}_{\mathcal{N}})\cup(\mathcal{O}\rightarrow\mathcal{C}_{\mathcal{N}})=\mathcal{C}_{\mathcal{N}}$, $(\mathcal{F}\rightarrow\mathcal{C}_{\mathcal{N}})\cap(\mathcal{O}\rightarrow\mathcal{C}_{\mathcal{N}})=\emptyset$.
For convenience, we denote all grids outside of $\mathcal{C}_{\mathcal{N}}$ as unpassable. If a path passes no unpassable grids, we denote it as collision-free; otherwise, it collides with obstacles.
III-A3 Neighborhood of grid
Denote $\delta(g)$ as the neighborhood of grid $g$, which consists of the closest $2\mathcal{N}$ grids near $g$.
III-A4 Agents
Assuming there are $k$ agents $A=\{a_{1},a_{2},\ldots,a_{k}\}$ in the grid space $\mathcal{C}_{\mathcal{N}}$, where each agent always occupies a passable grid. Their starting grids are denoted by $\mathcal{S}=\{s_{1},s_{2},\ldots,s_{k}\}$, and their target grids are denoted by $\mathcal{T}=\{t_{1},t_{2},\ldots,t_{k}\}$.
Each agent has a unique starting grid and a targeting grid. Specifically, for all $i\in\{1,2,\ldots,k\}$ and $j\in\{1,2,\ldots,k\}$ with $i\neq j$, we have $s_{i}\neq s_{j}$, $t_{i}\neq t_{j}$, $s_{i}\neq t_{j}$, and $t_{i}\neq s_{j}$.
III-A5 Time stamp
Time is discretized into timesteps. At each timestep, every agent can either move to an adjacent grid or wait at its current grid.
III-A6 Path
A path $p_{i}$ for agent $a_{i}$ is a sequence of grids that are pairwise adjacent or identical (indicating a wait action), starting at the start grid $s_{i}$ and ending at the target grid $t_{i}$. The cost of a path is the number of time steps it takes, i.e., the size of the path sequence $|p_{i}|$. We assume that the agents remain at their targets forever after completing their paths.
III-A7 Conflict
We say that two agents have a conflict if they are at the same grid or exchange their positions at the same timestep.
III-A8 Solution
A solution of a MAPF instance is a set of conflict-free paths $\{p_{1},p_{2},\ldots,p_{k}\}$, one for each agent. Generally, an optimal solution is one with minimal sum of costs (SOC) $\sum_{i=1}^{k}|p_{i}|$ or makespan $\mathop{\mathrm{argmax}}\limits_{i\in\{1,2,\ldots,k\}}|p_{i}|$. In this manuscript, we consider the effects of decomposition on both SOC and makespan.
III-A9 Solvability
If there exists a solution for a MAPF instance, it is solvable. In the manuscript, we only consider the decomposition of solvable MAPF instances.
A MAPF method is considered complete if it can always find a solution for a solvable problem; otherwise, it is incomplete. For example, EECBS is complete while PBS is incomplete.
A method to determine whether a MAPF instance is solvable is whether every agent can find a collision-free path (i.e., pass no unpassable grids) from the grid space. It is noteworthy that passing through other agents’ starting or target grids is acceptable.
III-A10 Decomposition of MAPF instance
Considering that the cost of solving a MAPF instance grows nearly exponentially as the number of agents increases [li2022mapf], decomposing a MAPF instance and solving subproblems typically incurs lower costs compared to solving the raw instance.
To reduce the cost of solve MAPF instance, we proposed a new technique: splitting a MAPF instance with multiple agents into multiple subproblems, each with fewer agents but under the same map, and solving these subproblems in a specific order. Here, we define the number of agents in each subproblem as the size of the subproblem. When solving a subproblem, it is crucial to ensure that the current paths do not conflict with the paths from previous subproblems. If every subproblem is solvable, combining all subproblems’ results yields the solution of the original MAPF instance.
The criteria for judging whether one decomposition is better than another involves sorting the subproblems’ sizes in decreasing order and comparing the sizes of the subproblems from the largest to the smallest. The first smaller subproblem encountered indicates the better decomposition.
For example, suppose there are two decompositions of a MAPF instance with 100 agents. The sizes of the subproblems for the first decomposition are 40, 20, 15, 14, and 11, while for the second decomposition, they are 40, 20, 19, 13, and 8. The first decomposition is considered better than the second one because the third cluster’s size (15) in the first decomposition is smaller than the second decomposition’s third subproblem size (19).
We compare the quality of decomposition in this manner because the cost of solving a MAPF instance grows exponentially as the number of agents increases [li2022mapf]. In theory, solving all agents together has a higher chance of finding a shorter solution compared to solving them separately. Therefore, MAPF methods with decomposition of MAPF instances tend to produce worse solutions than raw methods. However, we provide no theoretical analysis regarding the extent of this degradation in this manuscript; hence, we conduct empirical comparisons through extensive testing in Section VI.
In implementation, we measure the quality of decomposition by calculating the ratio between the size of the maximum subproblem and the number of agents of the raw MAPF instance. Since the cost of solving a MAPF instance is mainly determined by the maximum subproblem’s size, we refer to such a ratio as the decomposition rate in the following text.
Based on the criteria for comparing decompositions, a MAPF instance has a certain optimal decomposition that is better than all other decompositions. We have not found a direct way to obtain the optimal decomposition, but we can optimize an initial decomposition progressively to achieve a better decomposition.
III-B Theory basis of decomposition
A decomposition of a MAPF instance is considered legal if each subproblem remains solvable. Before discussing how to decompose a MAPF instance, we introduce how to check whether a decomposition of a MAPF instance is legal, i.e., every subproblem should have a solution that avoids conflicts with other subproblems’ paths.
We can simplify this by having each subproblem’s agents move until all previous subproblems have reached their targets, while avoiding the starts of subproblems that haven’t been solved yet. In this simplified scenario, every subproblem only needs to avoid previous subproblems’ targets and later subproblems’ starts. If we can find a solution for a subproblem under this simplified scenario, it is solvable. If we can find such a solution for every subproblem of a decomposition under the simplified scenario, the decomposition does not sacrifice the completeness of MAPF.
We check the solvability of a decomposition by verifying whether every agent of every subproblem can find a path under the simplified scenario. Essentially, the simplified scenario limits each subproblem’s agents to move only within a specific layer of the time-spatial space, where each layer has a unique timestep range. Therefore, we call our method LayeredMAPF. However, it is crucial to note that the solution under the simplified scenario may not be the only conflict-free solution, and the simplified scenario may generate solutions with extra wait actions. Thus, we will set the previous subproblem’s solution as constraints to avoid if a MAPF method supports it (serial MAPF methods). For MAPF methods that do not support setting external paths as constraints (parallel MAPF methods), we will set the previous subproblems’ targets and the later subproblems’ starts as occupied to ensure completeness.
To reduce the time cost of searching for paths, we propose a connectivity graph $\mathcal{G}_{c}$, whose nodes consist of free grid groups and the starts and targets of agents, and edges represent connectivity between nodes.
Definition 1
A free grid group is defined as a set of passable grids that contains no start or target grid of agents, and there always exists a path connecting any two grids within the group (where each waypoint in the path is within the same free grid group). We denote a free grid group as $f_{g}$ and a set of free grid groups as $F_{g}$. An example about free grid group are shown in Fig. 2.
A: a MAPF intance
B: free grid groups
C: connectivity graph
Figure 2: These figures depict a grid map and an MAPF instance containing three agents (A), free grid groups (B), and the corresponding connectivity graph. In these figures, ”0s, 1s, 2s” denote the starting positions of the three agents, while ”0t, 1t, 2t” represent their targets. Grey grids without labels indicate unpassable grids, while ”0f, 1f” label the grids belonging to two free grid groups, respectively.
Definition 2
The connectivity graph $\mathcal{G}_{c}$ is an undirected graph whose nodes consist of free grid groups, starting points, and targets for all agents, with edges representing the connectivity between these nodes.
Definition 3
The dependence path of an agent $a_{i}$ is defined as the path from $s_{i}$ to $t_{i}$, consisting of nodes in $\mathcal{G}_{c}$.
As an example, the dependence path for agent 0 in the instance shown in Fig. 2 is ”0s $\rightarrow$ 0f $\rightarrow$ 1t $\rightarrow$ 1f $\rightarrow$ 0t”; for agent 1, it is ”1s $\rightarrow$ 1f $\rightarrow$ 1t”; and for agent 2, it is ”2s $\rightarrow$ 1f $\rightarrow$ 2t”.
The dependence path of an agent reflects its relationship with other agents. For a valid decomposition, each agent’s path must not intersect with the targets of previous subproblems or the starts of later subproblems. The dependence paths of agents determine whether two agents must be solved in the same subproblem or the order in which they are solved across different subproblems. Specifically, there are four cases:
1, if the dependence path of agent $a_{i}$ contains only the start of agent $a_{j}$, it indicates that $a_{i}$ should be solved after $a_{j}$;
2, if the dependence path of agent $a_{i}$ contains only the target of agent $a_{j}$, it implies that $a_{i}$ should be solved before $a_{j}$;
3, if the dependence path of agent $a_{i}$ contains both the start and target of agent $a_{j}$, it signifies that $a_{i}$ and $a_{j}$ must be solved in the same subproblem;
4, if the dependence path of agent $a_{i}$ contains neither the start nor the target of agent $a_{j}$, it suggests that $a_{i}$ and $a_{j}$ can be solved in different subproblems, and the order of solving them does not matter.
It is worth noting that an agent’s dependence path is not unique. By updating their dependence paths, we can alter the relationship between two agents (in implementation, we check for the existence of such paths rather than explicitly maintaining all agents’ dependence paths). By changing the relationship between two agents, we can update the subproblems of a Multi-Agent Path Finding (MAPF) instance’s decomposition. Although there is no direct method to ensure that agents’ dependence paths correspond to smaller subproblems, we have proposed several approaches to encourage the generation of smaller subproblems; their details can be found in Section IV.
Here, we define the function for searching dependence paths as $p_{i}$ = $search\_path$($a_{i}$, $\mathcal{G}_{c}$, $available\_nodes$, $avoid\_nodes$). Here, $\mathcal{G}_{c}$ represents the connectivity graph of the MAPF instance, available nodes are those nodes in $\mathcal{G}_{c}$ that can be traversed, and avoid nodes are those nodes that must be avoided.
Specifically, an empty set for available nodes means that all nodes are available. If the $search\_path$ function succeeds, it returns a path; otherwise, it returns an empty set. If not specified, both available nodes and avoid nodes default to being empty.
In detail, the $search\_path$ function is implemented using a classic Best First Search approach, similar to A*. For this function, the cost and heuristic value represent the number of nodes the path has visited and the number of agents the path will visit. We consider a path to have visited an agent if it contains the start or target of that agent.
IV Methodology
As mentioned previously, we have not identified a direct method to decompose a MAPF instance into multiple subproblems. Thus, we initiate the process with an initial decomposition, progressively refining it to generate smaller subproblems while ensuring that each step maintains solvability. We define two types of subproblems encountered in various stages: clusters, which have no restrictions on the order of solving, and levels, which impose limitations on the order of solving.
The decomposition of a MAPF instance involves three steps:
1, identifying all agent dependencies and establishing initial clusters based on them;
2, refining agent dependencies to bipartition clusters until they cannot be further divided into smaller clusters;
3, further refining agent dependencies to decompose clusters into levels and sorting;
Step 1 and 2 are described in Section IV-A, and step 3 is described in Section IV-C.
An overview of the decomposition process of a MAPF instance is depicted in Figure 1.
Following decomposition, we explore methods to solve subproblems independently and then combine their results to obtain a conflict-free solution for the original MAPF instance, as detailed in Section IV-D.
IV-A Decomposetion to clusters
The first step is to determine the connectivity graph $\mathcal{G}_{c}$ from a grid space $\mathcal{C}_{\mathcal{N}}$ and the related agents $\mathcal{A}$. We then determine initial clusters based on the dependence paths derived from the connectivity graph.
Definition 4
Relevance of two agents: If an agent $a_{i}$’s dependence path contains another agent $a_{j}$’s start or target, we consider those two agents as relevant, regardless of whether $a_{j}$’s dependence path contains $a_{i}$’s start or target.
It is noteworthy that the relevance of two agents is determined by their dependence paths, so whether two agents are relevant changes as their dependence paths change.
For the three dependence paths mentioned before, agent 0 and agent 1 are relevant, while agent 2 is irrelevant to either of them.
Definition 5
Graph of agents’ relevance $\mathcal{G}_{a}$: an undirected graph where nodes represent agents and edges represent whether two agents are relevant.
Definition 6
Cluster: a maximal connected subgraph (i.e., connected component) of $\mathcal{G}_{a}$.
Intuitively, a cluster is a set of agents from the raw MAPF instance, where any agents within it are only relevant to agents within the same set (based on their dependence paths). A cluster is denoted as $r$, and a set of clusters is denoted as $\mathcal{R}$. Additionally, we refer to an agent set that constitutes a cluster as an independent agent set.
For the three dependence paths mentioned before, agent 0 and agent 1 form one cluster, while agent 2 forms another cluster.
An intuitive way to obtain clusters is by extracting clusters from the dependence paths that visit the fewest number of agents for each agent. We refer to such clusters as initial clusters.
However, it is evident that selecting dependence paths containing the fewest agents may not always result in the smallest possible clusters, as it does not consider how to guide dependence paths to avoid agents from other clusters.
There is room for decomposing these initial clusters into smaller ones by updating the dependence paths of agents (in implementation, we verify the existence of such paths rather than explicitly maintaining all agents’ dependence paths). Thus, we propose a method to iteratively bipartition initial clusters until further subdivision is not possible, aiming to minimize the size of subproblems. More details about bipartitioning can be found in the following section. This process ensures the solvability of decomposition, and the more bipartitioning steps, the better the results obtained, although it does not guarantee the discovery of the optimal decomposition. An overview of the process of decomposing an instance into clusters is outlined in Algorithm 1.
Input: $\mathcal{C}_{\mathcal{N}},\mathcal{G}_{c}$
Output: $\mathcal{R}$
1 ///// / determine initial clusters
2 $P=\{p_{1},p_{2},...,p_{k}\}$; ///// / dependence path of all agents
3 for i = 1, 2,…,$k$ do
4       $p_{i}=search\_path(a_{i},\mathcal{G}_{c},\emptyset,\emptyset,\emptyset)$;
5
6get $\mathcal{G}_{a}$ from $P$;
7 ///// / get initial clusters
8 $\mathcal{R}_{init}$ = all connected components of $\mathcal{G}_{a}$;
9 $\mathcal{R}=\emptyset$; ///// / final clusters
10 for $r\in\mathcal{R}_{init}$  do
11       $temp\_cluster=r$;
12       while $temp\_cluster\neq\emptyset$ do
13             $\{r_{major},r_{remain}\}$ = Bipartition of cluster ($r,\mathcal{G}_{c}$);
14             add $r_{major}$ to $\mathcal{R}$;
15             $temp\_cluster=r_{remain}$;
16
17
return $\mathcal{R}$;
Algorithm 1 Decomposition instance to clusters
IV-B Bipartition of clusters
This section is dedicated to decomposing a cluster into two smaller clusters. However, before delving into the decomposition process, it’s essential to introduce some necessary concepts.
Input: $r,\mathcal{G}_{c}$
Output: $\{r_{major},r_{remain}\}$
1 ///// / step 1
2 $\mathcal{G}_{u}=\emptyset$;
3 for $a_{i}\in r$ do
4       for $a_{j}\in\$r$ do
5             if $a_{i}\neq a_{j}$ then
6                   if $search\_path(a_{i},\mathcal{G}_{c},r,{a_{j}},\emptyset)=\emptyset$  then
7                         add $(a_{i}\rightarrow a_{j})$ as an edge to $\mathcal{G}_{u}$;
8
9
10
11
[…]
13 $r_{remain}=r-r_{major}$;
14 while $True$ do
15       ///// / step 2
16       $new\_agents=\emptyset$;
17       while $True$ do
18             for $a_{i}\in r_{remain}$ do
19                   if $search\_path(a_{i},\mathcal{G}_{c},r,r_{remain},\emptyset)=\emptyset$ then
20                         move $a_{i}$ from $r_{remain}$ to $r_{major}$;
21                         add $a_{i}$ to $new\_agents$;
22
23
24            if $r_{remain}$ is independent $or$ $r_{remain}$ is empty then
25                   break;
26
27      ///// / step 3
28       while $True$ do
29             for $a_{j}\in new\_agents$ do
30                   if $search\_path(a_{i},\mathcal{G}_{c},r_{major},r_{remain})=\emptyset$ then
31                         ///// / consider $a_{i}$ belong to cluster $r$, following search path always success $p_{i}=search\_path(a_{i},\mathcal{G}_{c},r,\emptyset,r\_major)$;
32                         move agents in ($p_{i}\cap r_{remain}$) from $r_{remain}$ to $r_{major}$;
33
34
35            if $r_{major}$ is independent then
36                   break;
37
38      ///// / step 4
39       if both $r_{major}$ and $r_{remain}$ are independent then
40             break;
41
return $\{r_{major},r_{remain}\}$;
Algorithm 2 Bipartition of cluster
Definition 7
A crucial concept to introduce is the notion of ”unavoidable agents” for an agent $a$ within a cluster $r$. Unavoidable agents of an agent $a$ in a cluster $r$ refer to those agents within $r$ that must be traversed by $a$’s dependence path. In other words, these agents must belong to the same cluster and cannot be further divided.
Definition 8
Unavoidance graph $\mathcal{G}_{u}$ of a cluster: an undirected representation of whether one agent is unavoidable to another agent within the cluster. It is important to note that for a given cluster $r$, its unavoidance graph is unique.
Both the unavoidance graph and the relevance graph depict relationships between agents within the cluster.
Definition 9
Maximum unavoidable agents of cluster: the largest connected component of the unavoidance graph $\mathcal{G}_{u}$ associated with the cluster.
Intuitively, these maximum unavoidable agents represent the largest undividable subset within the current cluster, are referred to as ”major set” during the cluster bipartition process. Agents within the cluster, excluding those belonging to the maximum unavoidable agents, are referred to as ”remaining set” during the cluster bipartition process.
It is noteworthy that during the bipartition process, agents in the remaining set will be moved to the major set.
Input: $r,\mathcal{G}_{c}$
Output: $\mathcal{L}$
1 ///// / determine initial dependence paths
2 $P=\emptyset$; ///// / $P[a]$ means agent a’s denpendence path
3 for agent $a$ in $r$ do
4       $p_{i}=search\_path(a_{i},\mathcal{G}_{c},r,{a_{j}})$;
5       add $p_{i}$ to $P$;
6
7get $\mathcal{G}_{s}$ from $P$;
8 determine all loop in $\mathcal{G}_{s}$ via strong component detect;
9 ///// / sort levels;
10 construct graph $\mathcal{G}_{l}$ via $\mathcal{G}_{s}$;
11 $root\_levels$ = level in $\mathcal{G}_{l}$ that not later than other level;
12 set order of all level to 0;
13 $levels$ = $root\_levels$;
14 while levels is not empty do
15       $next\_levels$ = $\emptyset$;
16       for level in levels do
17             for another_level that later than level in $\mathcal{G}_{l}$ do
18                   if $another\_level$’s order $\leq$ level’s order  then
19                         $another\_level$’s order = level’s order + 1;
20                         add $another\_level$ to $next\_levels$;
21
22
23
24      $levels$ = $next\_levels$;
25return all levels in the order of last visit to them;
Algorithm 3 Decomposing to levels and sorting
The bipartition of a cluster comprises four steps:
1, identify the maximum unavoidable agents of the cluster, referred to as the major set;
2, examine each agent in the remaining set to determine if its dependence path must pass through agents of the major set. Move agents meeting this criterion to the major set;
3, Check any newly added agents to the major set to ascertain if their dependence paths must pass through agents in the remaining set. Transfer such agents from agents the remaining set to the major set;
4, verify whether the major set and the remaining set both meet the requirements of the cluster. If they do (or if the remaining set is empty), exit and return the major set and the remaining set as the result of the bipartition. Otherwise, proceed to step 2;
The bipartition process of a cluster concludes when both sets satisfy the cluster’s requirements, or when there are no remaining agents, indicating that the cluster cannot be further decomposed into smaller clusters.
The pseudocode for bipartitioning a cluster is outlined in Algorithm 2.
Input: $\{P_{1},...,P_{m}\}$ ///// / assume there are $m$ subproblems and their solution is $P_{i}$
Output: $\{P_{1},...,P_{m}\}$
1 set last occupied time of all grids to 0;
2 for each solution $P$ in $\{P_{1},...,P_{m}\}$ do
3       while True do
4             $t$=1; ///// / current time index
5             ///// / check whether insert wait action to $P$
6             $need\_wait$ = False;
7             $all\_finished$ = True;
8             for path $p$ in $P$ do
9                   if t $>$ size of $p-1$ then
10                         ///// / don’t check path that have reach target
11                         continue;
12
13                  else
14                         $all\_finished$ = False;
15
16                  ///// / $p[t]$ means path $p$’s location at time $t$
17                   if last occupied time of $p[t]\geq t$ then
18                         $need\_wait$ = True;
19                         break;
20
21
22            if all_finished then
23                   break;
24            ///// / insert wait action to all path in $P$
25             if need_wait then
26                   for path $p$ in $P$ do
27                         if t $>$ size of $p-1$ then
28                               continue;
29
30                        repeat $p[t]$ in $p$ at time index $t$;
31
32
33            else
34                   ///// / update last occupied time of grids
35                   for path $p$ in $P$ do
36                         if t $>$ size of $p-1$ then
37                               continue;
38
39                        set last occupied time of $p[t]$ to $t$;
40
41
42            $t$ = $t$ + 1;
43
44
45return $\{P_{1},...,P_{m}\}$;
Algorithm 4 Merge results
Input: $\mathcal{C}_{\mathcal{N}},A$, MAPF
Output: $P$
1 ///// / 1, decomposition of MAPF instance
2 construct $\mathcal{G}_{c}$ from $\mathcal{C}_{\mathcal{N}},A$;
3 $R$ = Decomposing to cluster($\mathcal{G}_{c}$);
4 $L=\emptyset$;
5 for cluster $r$ in $R$ do
6       $L^{\prime}$ = Decomposing to leve and sorting($r$, $\mathcal{G}_{c}$);
7       merge $L^{\prime}$ to $L$;
8
[…]
10 ///// / 2, solve subproblems
11 for $l$ in $L$ do
12       if MAPF is serial then
13             set later subproblem’s start in $\mathcal{C}_{\mathcal{N}}$ to unpassable;
14             $current\_solution$ = MAPF($\mathcal{C}_{\mathcal{N}},l,all\_subproblem\_paths$);
15             add $current\_solution$ to $all\_subproblem\_paths$;
16             reset $\mathcal{C}_{\mathcal{N}}$;
17
18      else
19             set previous subproblem’s target and later subproblem’s start in $\mathcal{C}_{\mathcal{N}}$ to unpassable;
20             $current\_solution$ = MAPF($\mathcal{C}_{\mathcal{N}},l$);
21             add $current\_solution$ to $all\_subproblem\_paths$;
22             reset $\mathcal{C}_{\mathcal{N}}$;
23
24
25///// / 3, merge solutions of subproblems
26 if MAPF is serial then
27       $P$ = $all\_subproblem\_paths$;
28
29else
30       $P$ = Merge results($all\_subproblem\_paths$);
31
32return $P$;
Algorithm 5 Layered MAPF
IV-C Decomposing to levels and sorting
Clusters can be solved irrespective of the order of solving, thus providing an opportunity to decompose MAPF into smaller problems by considering the order of solving. By considering the order of solving, we can divide two agents into different subproblems even if one agent’s dependence path contains the start or target of another agent. We refer to these smaller problems decomposed from clusters as levels. In this section, we focus on how to decompose clusters into levels and determine the order of solving. To facilitate this discussion, we introduce some new concepts.
Definition 10
Order of solving agents: If an agent $a_{i}$ must be solved before another agent $a_{j}$, we denote it as $a_{i}>a_{j}$; if $a_{i}$ must be solved later than $a_{j}$, we denote it as $a_{i}<a_{j}$.
It is important to note that both $a_{i}>a_{j}$ and $a_{i}<a_{j}$ can coexist (indicating that both agents must be solved in the same problem) or neither can exist (indicating no order limitation between solving $a_{i}$ and $a_{j}$).
The order of solving agents is determined by whether an agent’s dependence path contains another agent’s start or target. If $a_{i}$’s dependence path contains $a_{j}$’s start, it implies that $a_{i}$ must be solved later than $a_{j}$ to ensure that $a_{j}$’s start is not occupied. Conversely, if $a_{i}$’s dependence path contains $a_{j}$’s target, $a_{i}$ must be solved before $a_{j}$ to ensure that $a_{j}$’s target is not occupied. Furthermore, the order of solving levels is determined by the order of solving agents within them.
Definition 11
Solving order graph $\mathcal{G}_{s}$ of a cluster: a directed graph representing whether an agent must be solved before another agent. Its nodes are agents, and edges indicate the order of solving between agents. The structure of the solving order graph is determined by each agent’s dependence path. Specifically, if agent $a_{i}$ must be solved before agent $a_{j}$, the corresponding edge in $\mathcal{G}_{s}$ is denoted as $a_{i}\rightarrow a_{j}$.
Definition 12
Level: a level is a strongly connected component of $\mathcal{G}_{s}$.
Intuitively, a level is a group of agents that form a loop in $\mathcal{G}_{s}$, indicating that the agents within the loop must be solved simultaneously. For example, if three agents $A$, $B$, and $C$ are in the same loop such that $A\rightarrow B\rightarrow C\rightarrow A$, it implies that $A$ must be solved before $B$, $B$ before $C$, and $C$ before $A$. Thus, $A$, $B$, and $C$ can only be solved simultaneously. Unlike clusters, where the order of solving is arbitrary, the order of solving levels is determined.
The order of solve levels is determine by the edge in $\mathcal{G}_{s}$ that connect them.
The order of solving levels is determined by the edges in $\mathcal{G}_{s}$ that connect them.
Definition 13
Order of solving levels: if a level $l_{i}$ must be solved before another level $l_{j}$, we denote it as $l_{i}>l_{j}$; if $l_{i}$ must be solved later than $l_{j}$, we denote it as $l_{i}<l_{j}$.
Definition 14
Level ordering graph $\mathcal{G}_{l}$ is a directed graph whose nodes are levels, and edges indicate whether a level must be solved earlier than another level. If there is no edge connecting two levels, this implies that there is no explicit order to solve them, although there may be an implicit order. For example, if level $A$ has no edge connecting to level $C$, but level $A$ is connected to level $B$ and level $B$ is connected to level $C$, it implies that level $A$ must be solved earlier than level $C$.
Essentially, $\mathcal{G}_{l}$ serves as a condensed version of $\mathcal{G}_{s}$.
There are five steps involved in decomposing a cluster into multiple levels and determining the order of solving:
1, determine each agent’s dependence path, and available agents constrained to the agents of the current cluster;
2, obtain the solving order graph $\mathcal{G}_{s}$ from the dependence paths of the current cluster;
3, identify all strong components of $\mathcal{G}_{s}$ as levels;
4, determine the relationships between levels by examining the edges that connect them. Construct the level ordering graph $\mathcal{G}_{l}$, which represents the order of solving between levels.
5, set levels that are not later than any other levels as root levels. Then, using Breadth First Search, traverse all levels in $\mathcal{G}_{l}$, where the order of the last visit to each level represents the order of solving them;
In implementation, Tarjan’s algorithm [tarjan1972depth] is utilized to determine strongly connected components from a directed graph.
It is important to note that after completing all steps of decomposition, we ensure that the final subproblems do not compromise solvability. The entire process of decomposing into levels and sorting them is depicted in Algorithm 3.
However, we cannot guarantee that the final subproblem is the optimal decomposition. Analogous to using gradient descent to solve non-linear optimization problems, we progressively get smaller and smaller subproblems, but the final result may still not be the global optimal solution.
IV-D Solving and combining
After completing the decomposition of a MAPF instance, the next step involves considering how to solve the subproblems separately and combine their results without conflicts. There are two main types of MAPF methods: serial MAPF methods and parallel MAPF methods. Serial MAPF methods, such as CBS-based methods, can use external paths as constraints to avoid conflicts, as they search for an agent’s path while the paths of other agents remain static. On the other hand, parallel MAPF methods, such as LaCAM and PIBT, cannot always treat external paths as dynamic obstacles to avoid, as they may avoid the same state (all agents’ locations) occur multiple times. For simplicity in this article, we do not set external paths as dynamic obstacles to avoid for all parallel MAPF methods.
In serial MAPF methods, we only need to merge separate results into a single set. However, for the second type of MAPF method, we need to add wait actions in the resulting paths to avoid conflicts between paths from different subproblems, as shown in Algorithm 4. In the following, we define serial MAPF as MAPF$(\mathcal{C}_{\mathcal{N}},\mathcal{A},P)$, and parallel MAPF methods as MAPF$(\mathcal{C}_{\mathcal{N}},\mathcal{A})$, where $\mathcal{A}$ represents agents and $P$ represents external paths that need to be avoided.
The overall process of decomposing a MAPF instance into multiple subproblems, solving the subproblems, and merging their results is illustrated in Algorithm 5.
As mentioned earlier, our decomposition method ensures the solvability of subproblems by ensuring the existence of a solution under the simplified scenario. However, our method does not guarantee that we find the optimal decomposition. While we have not provided theoretical analysis regarding this issue in this manuscript, we offer empirical analysis based on extensive testing, as detailed in Section VI.
V Results of decomposition
In theory, the decomposition of a MAPF instance reduces the cost of solving the instance, but it also incurs its own time and memory costs. Therefore, before analyzing how the decomposition of a MAPF instance contributes to solving the instance, it is essential to examine the time and memory usage of the decomposition process.
Specifically, we measure the peak memory usage as the memory usage, and record the memory usage of the program at every millisecond during implementation. However, due to the resolution of memory recording being 1KB, memory usage less than 1 KB is recorded as 0 MB. This phenomenon may occur under certain maps.
Furthermore, we analyze how each step of the decomposition process influences the overall decomposition by evaluating changes in the decomposition rate (maximum subproblem size / total number of agents) and the number of subproblems after each step. The steps are numbered sequentially, with the first step marked as ”1” and subsequent steps indicated as ”1, 2, 3,” and so forth.
We employ a classic MAPF dataset [stern2019mapf] as MAPF instances for our analysis, comprising 24 typical maps from the dataset. These maps feature an increasing number of agents, and each number of agents is randomly selected 100 times, resulting in a total of 22,300 MAPF instances. The same instances are utilized to evaluate how decomposition influences typical MAPF methods.
Experiments were conducted on a laptop running Ubuntu 20.04, equipped with a Ryzen 7 5800h (3.2GHz) CPU and 16GB of memory. All code is implemented in C++.
The results of our experiments are presented in Fig. 3 and Fig. 4.
1.empty-16-16
16*16 (256)
2.empty-32-32
32*32 (1,024)
3.maze-32-32-2
32*32 (666)
4.maze-32-32-4
32*32 (790)
5.maze-128-128-2
128x128 (10,858)
6.maze-128-128
-10
128x128 (14,818)
decomposition rate
subproblems
time cost (ms)
memory usage (MB)
7.den312d
65*85 (2,445)
8.den520d
256*257 (28,178)
9.Berlin_1_256
256*256 (47,540)
10.Paris_1_256
256*256 (47,240)
11.ht_chantry
162*141 (7,461)
12.lak303d
194*194 (14,784)
decomposition rate
subproblems
time cost (ms)
memory usage (MB)
Figure 3: These figures show our decomposition’s performance under various of maps, interms of decomposition rate, number of subproblems time cost and memory usage. We also present scale and number of free grids (in brackets).
13.random-32-32
-20
32x32 (819)
14.random-64-64
-10
64x64 (64x64)
15.room-32-32-4
32x32 (682)
16.room-64-64-8
64x64 (3,232)
17.room-64-64-16
64x64 (3,646)
18.warehouse-10-
20-10-2-1
161x63 (5,699)
decomposition rate
subproblems
time cost (ms)
memory usage (MB)
19.warehouse-10-
20-10-2-2
170x84 (9,776)
20.warehouse-20-
40-10-2-1
321x123 (22,599)
21.warehouse-20-
40-10-2-2
340x164 (38,756)
22.Boston_0_256
256x256 (47,768)
23.lt_gallowstemp
lar_n
251x180 (10,021)
24.ost003d
194x194 (13,214)
decomposition rate
subproblems
time cost (ms)
memory usage (MB)
Figure 4: These figures illustrate the performance of our decomposition method across various maps, focusing on decomposition rate, number of subproblems, time cost, and memory usage. Additionally, we provide information on the scale and number of free grids (in brackets).
V-A Decomposition rate
The decomposition rate is closely correlated with the time cost and memory usage in solving a MAPF instance, as both time cost and memory usage are primarily determined by the size of the largest subproblem.
V-A1 Performance under different maps
As depicted in Fig. 3, the final decomposition rate (i.e., the decomposition rate after step 3) demonstrates a consistent increase across almost all maps, with the exception of maps containing a high number of free grids (e.g., den520, Berlin_1_256). In these maps, the number of free grids closely aligns with the number of agents, resulting in fewer opportunities to alter dependence paths and decompose the instance. Consequently, the size of the largest subproblem approaches that of the original MAPF instance. This limitation becomes increasingly significant as the number of agents grows, ultimately leading to a decomposition rate close to 1, indicating that decomposition becomes ineffective when agents are densely packed.
Conversely, in maps with a surplus of free grids compared to the number of agents, such as den520 and Berlin_1_256, decomposition remains highly effective despite increases in the number of agents. In these scenarios, the abundance of free grids results in numerous free grid groups, facilitating the generation of dependence paths that traverse multiple agents’ start and target locations.
V-A2 How each step contribute to decomposition
As shown in Fig. 3, we observe that each step of the decomposition process contributes to a decrease in the decomposition rate (i.e., a reduction in the maximum size of subproblems) overall. However, the extent of their contributions varies across different maps.
In maps characterized by a surplus of free grids, exceeding the number of agents (such as Berlin_1_256, Paris_1_256, and Den520d), the maximum size of clusters is reduced to merely 1 after step 1, even with hundreds of agents. This is due to the high likelihood of an agent’s initial dependence path avoiding passing through other agents’ start or target locations. Consequently, subsequent steps 2 and 3 have no further room for decomposition, resulting in similar time costs after each step in these maps.
Conversely, in maps where the number of free grids is close to the number of agents (e.g., empty_16_16, empty_32_32, random-32-32-20, and maze-32-32-4), steps 2 and 3 contribute more to decomposition than step 1 as the number of agents increases. This is because there is a lower likelihood of an agent’s initial dependence path avoiding passing through other agents’ start or target locations. Consequently, steps 2 and 3 have the opportunity to decompose subproblems by updating dependence paths.
V-B Number of subproblems
The count of subproblems generally follows a pattern of initial increase followed by a decrease as the number of agents increases. Initially, decomposition easily generates small subproblems, leading to an increase in their count as the number of agents grows. However, as agents become denser, decomposing becomes more challenging, resulting in an increase in the decomposition rate and a subsequent decrease in the number of subproblems until it equals 1, indicating the raw MAPF instance.
Special cases arise in maps with numerous free grids, exceeding the number of agents, such as den520d and Berlin_1_256. In these instances, decomposition divides the raw instance into subproblems, each containing only one agent, resulting in the number of subproblems equaling the number of agents in the raw MAPF instance. However, it is predictable that as the number of agents increases, decomposition will become increasingly ineffective, ultimately reducing the number of subproblems to 1.
V-C Time cost and memory usage
Time cost and memory usage are crucial factors in the application of decomposition for MAPF. While theoretically, decomposing a MAPF instance should reduce the time cost and memory usage of solving the problem, if the decomposition process itself consumes too much time or memory space, it may not effectively reduce the total cost of solving the MAPF instance. So we analysis how many resources decomposition cost under various of maps.
Generally, as depicted in the Fig. 3 and Fig. 4, the time cost of decomposition is less than 1 second in most maps and less than 3 seconds in the worst-case scenario (such as 800 to 1000 dense agents), which is deemed acceptable for practical application. In the majority of cases, all three steps contribute to the time cost. However, there are differences in the time cost of steps across different maps. In maps with an abundance of free grids exceeding the number of agents, such as den512d, Berlin_1_256, and Paris_1_256, step 1 consumes almost all of the time cost, resulting in overlapping time costs after step 1, step 2, and step 3. Conversely, in maps with few free grids close to the number of agents, all steps contribute to the time cost, as these scenarios require steps 2 and 3 to decompose initial subproblems into smaller subproblems.
Regarding memory usage, decomposition for almost all maps consumes less than 1 MB, which is considered acceptable for common platforms such as laptops or industrial process computers.
V-D Summary
Our method demonstrates high effectiveness for MAPF instances with free grids exceeding two times the number of agents, while its effectiveness diminishes as the number of free grids approaches two times the number of agents. This is because a higher number of free grids increases the likelihood of decomposing the instance into smaller subproblems. Conversely, for the same map, the effectiveness of decomposition decreases with an increase in the number of agents in the MAPF instance.
Regarding subproblems, in most maps, the number of subproblems initially increases and then decreases as the number of agents increases. Initially, with only a few agents, decomposition is effective, but as the number of agents increases, decomposition becomes less effective.
In terms of costs, the average time cost is less than 1 second in average, with a maximum of less than 3 seconds in the worst cases (such as 800 to 1000 dense agents), and memory usage remains below 1 MB. Comparatively, maps with more free grids than the number of agents require fewer computations and less memory space.
VI Results of decomposition’s application
[…]
As a common rule in solving MAPF instances, we set an upper bound on the time cost (30 seconds). Methods that are complete but fail to find a solution within the limited time are considered as failed. Additionally, methods that run out of memory space are also considered as failed.
Figure 5: These figures provide a summary comparison between the raw MAPF method and the layered MAPF method, combining the results across all maps. In the visualizations, crossed bars represent the layered version of the MAPF method, while empty bars represent the raw MAPF method. The same applies hereinafter.
EECBS
PBS
LNS2
PushAndSwap
[…]
PIBT+
LaCAM
Figure 6: These figures serve as a detailed version of Fig. 5, offering insights into the performance of both raw MAPF methods and layered MAPF methods under specific maps.
VI-A Explicit Estimation CBS
Explicit Estimation CBS (EECBS) [li2021eecbs] stands out as one of the state-of-the-art CBS-based MAPF methods, characterized by its completeness and bounded suboptimality. Beyond its explicit estimation in high-level search, EECBS also incorporates conflict resolution techniques such as bypassing conflicts, prioritizing conflicts, symmetry reasoning, and weighted dependency graphs to reduce the time cost of CBS. Further details can be found in Section II. EECBS’s code  [^3] https://github.com/Jiaoyang-Li/EECBS operates by searching each agent’s path separately and provides interfaces to avoid conflicts with external paths, rendering it a serial MAPF method. In its layered version, which we refer to as ”layered EECBS,” previous subproblems’ paths and subsequent subproblems’ starts are set as constraints.
As illustrated in Fig. 5, on average, layered EECBS exhibits lower time costs ($1.0\times 10^{4}$ ms $<$ $1.5\times 10^{4}$ ms), leading to a higher success rate (0.77 $>$ 0.64). The time cost of decomposition is relatively small compared to the time cost of solving the MAPF instance. Regarding memory usage, layered EECBS consumes less memory space compared to raw EECBS ($12$ MB $<$ $24$ MB). Decomposition of the MAPF instance significantly reduces the growth of the conflict tree’s size by decreasing the number of agents solved simultaneously.
In terms of path quality, layered EECBS yields a larger makespan (layered EECBS: $2.5\times 10^{2}$, raw EECBS: $2.0\times 10^{2}$), resulting in a larger sum of costs (layered EECBS: $5.0\times 10^{4}$, raw EECBS: $3.4\times 10^{4}$). In maps where both raw EECBS and layered EECBS achieve a success rate close to 1 (e.g., map 1, 8, 10, 19, 21 and 22), both methods exhibit similar makespan and sum of costs. However, in maps where raw EECBS has a lower success rate compared to layered EECBS (e.g., map 4, 16, 17 and 24), layered EECBS demonstrates a larger makespan and sum of costs, as depicted in Fig. 6.
This phenomenon can be attributed to two factors: Firstly, solving all agents together provides more opportunities to find shorter solutions than solving them separately. Secondly, raw EECBS struggles to find longer solutions, particularly in maps where its success rate is lower.
In summary, layered EECBS boasts explicit advantages in terms of time cost, memory usage, and success rate compared to raw EECBS. Additionally, in maps where both methods achieve a high success rate, their solutions tend to be similar. However, layered EECBS outperforms raw EECBS in finding longer solutions when the latter fails to do so.
VI-B Priority Based Search
Priority Based Search (PBS) [ma2019searching] is an incomplete, suboptimal, and priority-based MAPF method. Conceptually, when faced with a collision, PBS greedily chooses which agent should be given higher priority. It backtracks efficiently and explores other branches only if there is no solution in the current branch. PBS searches each agent’s path separately and it’s code [^4] https://github.com/Jiaoyang-Li/PBS provides interfaces to avoid conflicts with external paths, making it a serial MAPF method. In its layered version, which we refer to as ”layered PBS,” we set previous subproblems’ paths and subsequent subproblems’ starts as constraints.
As shown in Fig. 5, on average, layered PBS has a lower time cost ($1.2\times 10^{4}$ ms $<$ $2.0\times 10^{4}$ ms) and thus a higher success rate (0.75 $>$ 0.42). The time cost of decomposition is relatively small compared to the time cost of solving the MAPF instance. In terms of memory usage, layered PBS has lower memory usage compared to raw PBS (1.0 MB $<$ 3.9 MB) because decomposition of the MAPF instance reduces the complexity of the priority tree significantly.
When it comes to path quality, layered PBS has a larger makespan ($2.5\times 10^{2}$ $>$ $1.3\times 10^{2}$) and thus a larger SOC ($5.5\times 10^{4}$ $>$ $1.7\times 10^{4}$). Similar to EECBS, raw PBS finds it difficult to find longer solutions and can only finds shorter solutions, resulting in lower success rates for maps where raw PBS struggles (e.g., map 4, 6, 16 and 17). For maps where both methods have high success rates (e.g., map 1, 9 and 10), they have similar makespan and SOC, as shown in Fig. 6.
In summary, layered PBS has explicit advantages in time cost, memory usage, and success rate compared to raw PBS. Additionally, layered PBS produces solutions similar to raw PBS’s solutions on maps where both methods have success rates close to 1. Furthermore, layered PBS achieves a higher success rate in finding long solutions when raw PBS fails.
VI-C Large Neighborhood Search 2
Large Neighborhood Search 2 (LNS2) [li2022mapf] is a complete and suboptimal MAPF method that starts from a set of paths with collisions and repeatedly replans subsets of paths to reduce the overall number of collisions until the paths become collision-free. LNS2 updates each agent’s path separately, and its code [^5] https://github.com/Jiaoyang-Li/MAPF-LNS provides interfaces to avoid conflicts with external paths, making it a serial MAPF method.
As shown in Fig. 5, on average, layered LNS2 has a lower time cost ($1.0\times 10^{4}$ ms $<$ $1.4\times 10^{4}$ ms) and thus a higher success rate (0.76 $>$ 0.55). In terms of memory usage, layered LNS2 has lower memory usage compared to raw LNS2 (2.4 MB $<$ 32 MB). Due to the decomposition of the MAPF instance, other agents’ paths are considered as dynamic obstacles and do not join conflict resolve, resulting in layered LNS2 requiring fewer repair actions to fix conflicts between agents’ paths.
Regarding path quality, layered LNS2 has a larger makespan ($2.5\times 10^{2}$ $>$ $2.1\times 10^{2}$) and SOC ($5.1\times 10^{4}$ $>$ $3.4\times 10^{4}$). Similar to EECBS and PBS, raw LNS2 struggles to find longer solutions and can only discovers shorter solutions, particularly in maps where raw LNS2 has a lower success rate (e.g., map 12 and 16). For maps where both methods have similar success rates (e.g., map 1, 2, 8, 21 and 22), raw LNS2 and layered LNS2 have similar makespan and SOC, as shown in Fig. 6.
In summary, layered LNS2 exhibits explicit advantages in time cost, memory usage, and success rate compared to raw LNS2. Additionally, layered LNS2 produces solutions similar to raw LNS2’s solutions in maps where both methods have success rates close to 1. Furthermore, layered LNS2 achieves a higher success rate in finding long solutions when raw LNS2 cannot.
VI-D Push and Swap
Push and Swap (PAS) [luna2011push] is a suboptimal and complete rule-based MAPF method. The algorithm employs two primitives: a “push” operation where agents move towards their goals until no progress can be made, and a “swap” operation that allows two agents to exchange positions without affecting other agents’ configurations.
For each agent in the MAPF instance, the algorithm uses “push” to move the agent toward its goal along the shortest path until encountering another agent. Then, it employs “swap” to exchange their positions without altering other agents’ positions. If a “swap” fails, the problem is deemed unsolvable. The algorithm alternates between ”push” and “swap” operations until all agents reach their targets.
Push and Swap updates agents’ paths simultaneously (via an action called “MULTIPUSH”, which moves a set of adjacent agents simultaneously), and its code [^6] https://github.com/Kei18/pibt2 provides no interfaces to avoid conflicts with external paths and ensures that the same state of agents does not occur twice in the solution, making it a parallel MAPF method. In its layered version, referred to as ”layered Push and Swap” or ”layered PAS,” we set previous subproblems’ targets and subsequent subproblems’ starts as static obstacles. Subsequently, we merge subproblems’ solutions by adding wait actions after solving all subproblems.
As shown in Fig. 5, on average, layered PAS consumes less time ($4.6\times 10^{3}$ ms $<$ $8.8\times 10^{3}$ ms) and memory space (6.2 MB $<$ 70 MB) than raw PAS, resulting in a higher success rate (0.95 $>$ 0.82). Decomposing the MAPF instance enables layered Push and Swap to solve subproblems separately, requiring fewer “push” and “swap” actions. However, the number of “push” and “swap” actions grows exponentially as the number of agents increases.
Regarding path quality, as PAS is a parallel MAPF method, the addition of wait actions to merge subproblems’ solutions, layered PAS has a larger makespan ($1.0\times 10^{4}$ $>$ $6.1\times 10^{3}$) and SOC ($3.6\times 10^{6}$ $>$ $2.4\times 10^{6}$) compared to raw PAS.
In summary, layered Push and Swap consume less time and memory space than raw Push and Swap, resulting in a higher success rate. However, the introduction of wait actions to merge subproblems’ solutions leads to a larger makespan and SOC compared to raw Push and Swap.
[…]
Hierarchical Cooperative A* (HCA*) [silver2005cooperative] is a suboptimal and incomplete prioritized MAPF algorithm that decouples into a series of single-agent searches. The individual searches are performed in three-dimensional space-time, taking into account the planned routes of other agents. Essentially, HCA* decomposes the MAPF instance into subproblems that involve only one agent. HCA* also calculates a distance table for each agent to determine their priority.
HCA* searches each agent’s path separately, making it a serial MAPF method. But the code [^7] https://github.com/Kei18/pibt2 for HCA* provides no interfaces to set external paths as dynamic obstacles to avoid. We updated its code to enable it to set external paths as dynamic obstacles.
On average, layered HCA* has a slightly higher time cost than raw HCA* ($1.3\times 10^{4}$ ms $>$ $1.1\times 10^{4}$ ms) and a similar success rate (layered HCA*: 0.68, raw HCA*: 0.64) as HCA* solves each agent separately, and decomposition does not improve the efficiency of HCA*. However, layered HCA* requires less memory space than raw HCA* ($2.9\times 10^{2}$ MB $<$ $5.4\times 10^{2}$ MB), as layered HCA* only stores the distance table of the current subproblem’s agents, while raw HCA* needs to store the distance table of all agents.
Regarding path quality, layered HCA* has a slightly larger makespan ($2.4\times 10^{2}$ $>$ $2.1\times 10^{2}$) and SOC ($4.6\times 10^{4}$ $>$ $3.8\times 10^{4}$). Because HCA* solves agents in order of decreasing distance to the target, raw HCA* solves agents with longer paths first, limiting the makespan of these agents’ solutions, while layered HCA* breaks this order as agents with long paths may not be in the same subproblem and may not be solved first.
In summary, layered HCA* shows little improvement compared to raw HCA* and has slightly worse solutions because HCA* has a fixed priority and solves each agent separately. However, layered HCA* has an advantage in memory usage, as it only needs to store the distance table of the current subproblem’s agents rather than all agents.
VI-F Priority Inheritance with Backtracking$+$
Priority Inheritance with Backtracking$+$ (PIBT$+$) [okumura2022priority] is an incomplete and suboptimal MAPF algorithm. Its high level assigns a unique priority to each agent at every timestep, while its low level moves each agent one step at every timestep. Consequently, all intermediate states are saved until all agents reach their targets. PIBT$+$ avoids the occurrence of the same state of agents twice in the solution. And the code [^8] https://github.com/Kei18/pibt2 for PIBT$+$ provides no interfaces to avoid conflicts with external paths, making it a parallel MAPF method.
In general, layered PIBT$+$ takes more time than raw PIBT$+$ ($2.0\times 10^{3}$ ms $>$ $6.9\times 10^{2}$ ms). As mentioned in Section V, the decomposition of MAPF instances with 800 to 1000 agents may cost 1s to 3s, and these instances play a major role in contributing to the mean time cost of the MAPF method. Although decomposition reduces the time cost of PIBT$+$, the time cost of decomposition offsets the reduced time cost and thus increases the overall time cost.
Layered PIBT$+$ has a similar success rate compared to raw PIBT$+$ (layered PIBT$+$: 0.99, raw PIBT$+$: 0.99), as both layered PIBT$+$ and raw PIBT$+$ typically take less than 30s in most cases. In terms of memory usage, layered PIBT$+$ requires less memory space than raw PIBT$+$ (6.8 MB $<$ 18 MB), as decomposition reduces the number of intermediate states by reducing the number of agents solved simultaneously.
Regarding solution quality, as PIBT$+$ is a parallel MAPF method, layered PIBT$+$ introduces extra wait actions, resulting in solutions with higher makespan ($9.5\times 10^{3}$ $>$ $4.3\times 10^{2}$) and SOC ($3.1\times 10^{6}$ $>$ $1.2\times 10^{5}$) compared to raw PIBT$+$’s solutions.
In summary, layered PIBT$+$ shows little improvement compared to raw PIBT$+$ and has worse solutions due to the introduction of wait actions when merging solutions. However, layered PIBT$+$ has an advantage in memory usage, as it reduces the number of intermediate states by reducing the number of agents solved simultaneously.
VI-G Lazy Constraints Addition for MAPF
Lazy Constraints Addition for MAPF (LaCAM) [okumura2023lacam] is a complete and suboptimal method. More details about LaCAM can be found in Section II. LaCAM avoids the occurrence of the same state of agents twice in the solution. The code [^9] https://github.com/Kei18/lacam for LaCAM provides no interfaces to avoid conflicts with external paths, making it a parallel MAPF method.
In general, layered LaCAM has slightly lower time cost than raw LaCAM ($3.6\times 10^{3}$ ms $<$ $3.8\times 10^{2}$ ms). Similar to PIBT$+$, the decomposition of MAPF instances with 800 to 1000 agents may cost 1s to 3s, and these instances play a major role in determining the time cost of the MAPF method. Although decomposition reduces the time cost of LaCAM, the time cost of decomposition offsets the reduced time cost, resulting in no significant difference in the overall time cost.
Layered LaCAM has a slightly better success rate compared to raw LaCAM (layered LaCAM: 0.93, raw LaCAM: 0.88), as both layered LaCAM and raw LaCAM typically take less than 30s in most cases. In terms of memory usage, layered LaCAM requires less memory space than raw LaCAM (12 MB $<$ 18 MB), as decomposition reduces the number of intermediate states by reducing the number of agents solved simultaneously.
Regarding solution quality, layered LaCAM introduces extra wait actions, resulting in solutions with higher makespan ($9.3\times 10^{3}$ $>$ $3.2\times 10^{2}$) and SOC ($3.1\times 10^{6}$ $>$ $8.6\times 10^{4}$) compared to raw LaCAM’s solutions.
In summary, layered LaCAM shows little improvement compared to raw LaCAM and has worse solutions due to the introduction of wait actions when merging solutions. However, layered LaCAM has an advantage in memory usage, as it reduces the number of intermediate states by reducing the number of agents solved simultaneously.
VI-H Summary
On average, decomposition of MAPF instances significantly reduces the memory usage of all seven methods and also decreases the time cost for EECBS, PBS, LNS2, Push and Swap. Layered MAPF methods exhibit higher success rates compared to raw MAPF methods, particularly for serial MAPF methods, with the exception of HCA*, an incomplete prioritized method that solves agents one-by-one.
Regarding solution quality, the layered version of serial MAPF methods generally exhibits similar quality compared to the raw version. However, the layered version of parallel MAPF methods tends to produce worse solutions due to the insertion of numerous wait actions during solution merging. This disparity arises from the fact that serial MAPF methods treat external paths as dynamic obstacles to avoid, while parallel MAPF methods do not.
In conclusion, decomposition of MAPF instances is most advantageous for serial MAPF methods, resulting in reduced time cost and memory usage without significantly sacrificing solution quality. For parallel MAPF methods, decomposition reduce memory usage significantly but often leads to worsened solutions without notable improvements in time cost.
Moreover, serial MAPF methods (e.g., EECBS, LNS2) generally demand more time compared to parallel MAPF methods (e.g., LaCAM and PIBT+), yet they yield higher-quality solutions.. Notably, serial MAPF methods frequently fail due to time constraints, while parallel MAPF methods tend to fail due to memory limitations.
The observed discrepancy can be attributed to the differing search algorithms utilized: serial methods repeatedly search for complete paths. However, they employ memory-efficient data structures such as priority trees or conflict trees to store knowledge about resolving conflicts. In contrast, parallel methods store previous conflict-solving knowledge as intermediate states to avoid searching for the complete path of agents repeatedly. However, they may exhaust memory resources under complex MAPF instances.
VII Conclusion
Motivated by the exponential growth in the cost of solving MAPF instances (in terms of time and memory usage) as the number of agents increases, we proposed layered MAPF as a solution to reduce the computational burden. This approach decomposes a MAPF instance into multiple smaller subproblems without compromising solvability. Each subproblem is solved in isolation, with consideration given to other subproblems’ solutions as dynamic obstacles. Our methodology involves a progressive decomposition of MAPF instances, ensuring that each step preserves solvability.
In the results of our decomposition of MAPF instances (Section V), we observed that our method is highly effective for MAPF instances with free grids exceeding twice the number of agents. On average, the time cost is around 1s and never exceeds 3s, even for dense instances with 800 to 1000 agents. Memory usage remains below 1MB, with fewer computations and memory space required for maps with more free grids than agents.
When applied to the state-of-the-art methods (EECBS, PBS, LNS, HCA*, Push and Swap, PIBT+, LaCAM), layered MAPF significantly reduces memory usage and time cost, particularly for serial MAPF methods. Consequently, layered MAPF methods achieve higher success rates than raw MAPF methods, especially for serial MAPF. And the quality of solution for the layered version of serial MAPF methods is similar to the raw version, while the layered version of parallel MAPF methods produces inferior solutions due to the introduction of numerous wait actions during solution merging.
In conclusion, decomposition of MAPF instances is most beneficial for serial MAPF methods, resulting in reduced time cost and memory usage without sacrificing solution quality significantly. However, for parallel MAPF methods, decomposition may reduce memory usage but often worsens the solution without notable improvements in time cost.
Despite its effectiveness, layered MAPF has limitations: it becomes less effective as the number of agents increases in dense instances, and its application to parallel MAPF methods introduces numerous wait actions during solution merging.
In the future, we plan to propose new merging solution techniques for parallel methods without compromising solution quality. Additionally, we aim to generalize the idea of decomposing MAPF instances to address extensions of MAPF problems, such as considering the shape of agents.
Acknowledgments
The first author thanks Lu Zhu for her encouragement and support during the consummation of Layered MAPF.
[…]